# HalluClean: A Unified Framework to Combat Hallucinations in LLMs

### v3

@@ Introduction
Large language models (LLMs) have revolutionized natural language processing (NLP), powering applications such as conversational agents, content generation, and decision support systems (Chowdhery et al. 2023; Touvron et al. 2023; Bang et al. 2023; Qin et al. 2023). These models leverage large-scale pretraining and are further enhanced via instruction tuning (Chung et al. 2024; Wang et al. 2022b, a) and alignment techniques that optimize for human preferences (Ouyang et al. 2022; Achiam et al. 2023). However, despite their remarkable fluency and versatility, LLMs frequently produce hallucinated or factually incorrect content (Huang et al. 2023; Ji et al. 2023), undermining their reliability in safety-critical contexts.
@@
Existing hallucination mitigation approaches typically fall into two categories. Retrieval-augmented generation methods (Varshney et al. 2023; Cao et al. 2023; Kang, Ni, and Yao 2023; Rawte et al. 2023) query external knowledge sources to validate or correct model outputs. Meanwhile, supervised detection approaches (Razumovskaia et al. 2024; Zhang et al. 2023; Qiu et al. 2023) rely on human-labeled data to train classifiers that identify hallucinations. While both strategies have shown promise, they suffer from key limitations: retrieval-based methods depend on the availability and accuracy of external sources, and annotation-based methods are costly and poorly generalize to novel hallucination types. Furthermore, hallucination behaviors vary widely across tasks—such as question answering (Zheng, Huang, and Chang 2023), summarization (Cao, Dong, and Cheung 2022), and dialogue (Das, Saha, and Srihari 2022)—yet most prior work focuses narrowly on specific settings (Mündler et al. 2023), limiting scalability and robustness.
@@
Hallucinations in large language models (LLMs) have been extensively studied, focusing on their causes (Pan et al. 2023; Chen and Shu 2024; Kasai et al. 2024; Wang et al. 2023a; Lee et al. 2022; Yao et al. 2023), evaluation methodologies (Lin, Hilton, and Evans 2021; Lee et al. 2022; Min et al. 2023; Li et al. 2023a), and behavioral analysis (Zhao et al. 2023; Dong et al. 2024; Li et al. 2023b). Many studies have investigated ways to mitigate hallucinations through retrieval-augmented generation (Peng et al. 2023; Varshney et al. 2023; Kang, Ni, and Yao 2023) and supervised fine-tuning (Elaraby et al. 2023; Razumovskaia et al. 2024; Zhang et al. 2023). To improve LLM reliability, researchers have explored prompting-based solutions. For example, Si et al. (2022) proposed simple yet effective prompts that enhance GPT-3’s factual accuracy, while Mitchell et al. (2022) introduced a two-model framework in which one model generates responses and another evaluates their logical coherence. More recently, Mündler et al. (2023) found that 17.7% of ChatGPT-generated sentences contain self-contradictions and proposed a three-step pipeline to detect and mitigate them without relying on external knowledge. Despite these advancements, hallucination detection and mitigation in LLMs remain challenging. Our method uses task-adaptive prompts in a zero-shot setting to guide LLMs in detecting and revising hallucinated content by eliciting and leveraging model reasoning.
@@
Prompting strategies have played a crucial role in enhancing the reasoning abilities of LLMs. Chain-of-Thought (CoT) prompting (Wei et al. 2022) explicitly structures intermediate reasoning steps, significantly improving model performance on complex reasoning tasks. Building upon this, various enhancements have been proposed, including prompt ensembling (Wang et al. 2022a; Li et al. 2022; Fu et al. 2022), problem decomposition (Zhou et al. 2022; Khot et al. 2022; Dua et al. 2022), and structured planning methods (Yao et al. 2022; Huang et al. 2022; Wang et al. 2023b; Liu et al. 2023). To reduce manual effort and computational overhead, zero-shot CoT prompting (Kojima et al. 2022) was introduced, allowing LLMs to autonomously generate reasoning steps without the need for labeled exemplars. However, existing methods primarily focus on general reasoning tasks, and limited work has explored their applicability in addressing hallucinations within LLM-generated text.
@@
Contradictions within the same response (e.g., mutually exclusive statements) signal hallucination. Such contradictions appear in  17.7% of ChatGPT-generated sentences (Mündler et al. 2023).
@@
The core innovation of HalluClean lies in its structural reasoning mechanism for hallucination detection. While direct classification can identify obvious hallucinations, we find that step-by-step reasoning significantly improves detection accuracy, especially for subtle or complex hallucination cases. Our approach draws inspiration from cognitive science literature on human reasoning, which emphasizes the role of structured thinking in error detection and verification (Kahneman 2011).
@@
We collect evaluation data from four established hallucination detection benchmarks: 1. HaluEval (Li et al. 2023a): Covers hallucinated samples across three task types—question answering, knowledge-grounded dialogue, and text summarization. 2. UMWP (Sun et al. 2024): Evaluates hallucination in math word problems (MWPs) by identifying questions with no or non-unique solutions. Such unanswerable questions are known to induce hallucinations in LLMs and are often used to test whether models can recognize ill-posed or unsolvable problems—similar to how educators gauge student understanding with trick questions. 3. ChatProtect (Mündler et al. 2023):Focuses on self-contradictory hallucinations, where a language model produces logically inconsistent statements within the same context. 4. HaluBench (Ravi et al. 2024):A domain-specific benchmark composed of hallucinated QA examples in the medical and financial domains, sourced from CovidQA, PubMedQA, and FinanceBench.
@@
The effectiveness of the proposed framework is evaluated along three dimensions: 1. Hallucination Reduction Rate: To measure the effectiveness of the revision step, we compute the hallucination reduction rate by comparing the number of hallucinations detected before and after revision. Specifically, we first identify hallucinations in the original outputs, then re-evaluate the revised outputs using GPT-4o-mini. 2. Revision Success Rate: To assess revision quality, we compute BERTScore between each revised output and its gold reference. A revision is considered acceptable if the BERTScore (Zhang et al. 2019) exceeds 0.85(chosen to balance strict semantic fidelity and flexibility in surface expression). The revision success rate is the proportion of acceptable revisions among all hallucinations identified before revision. For multi-word problems (MWPs), revision quality evaluation is performed based on unanswerable reason categories, using exact match between predicted and gold labels. 3. Hallucination Detection: Since hallucination detection is formulated as a binary classification task, we evaluate its effectiveness using two standard metrics: F1 score and accuracy, computed against human-annotated gold labels from the original benchmarks. Accuracy reflects overall correctness, while F1 provides additional insight into the model’s balance between precision and recall, ensuring that both metrics are measured with respect to verified ground truth.
@@
To evaluate whether our method complements external knowledge, we test hallucination detection on the QA task from HalEval (Li et al. 2023a) under two settings: vanilla (no external knowledge) and retrieval-augmented (with background knowledge).

### v4

@@ Introduction
Large language models (LLMs) have revolutionized natural language processing (NLP), powering applications such as conversational agents, content generation, and decision support systems (chowdhery2023palm; touvron2023llama; bang2023multitask; qin2023chatgpt). These models leverage large-scale pretraining and are further enhanced via instruction tuning (chung2024scaling; wang2022super; wang2022self) and alignment techniques that optimize for human preferences (ouyang2022training; achiam2023gpt). However, despite their remarkable fluency and versatility, LLMs frequently produce hallucinated or factually incorrect content (huang2023survey; ji2023survey), undermining their reliability in safety-critical contexts.
@@
Existing hallucination mitigation approaches typically fall into two categories. Retrieval-augmented generation methods (varshney2023stitch; cao2023step; kang2023ever; rawte2023troubling) query external knowledge sources to validate or correct model outputs. Meanwhile, supervised detection approaches (razumovskaia2024dial; zhang2023r; qiu2023think) rely on human-labeled data to train classifiers that identify hallucinations. While both strategies have shown promise, they suffer from key limitations: retrieval-based methods depend on the availability and accuracy of external sources, and annotation-based methods are costly and poorly generalize to novel hallucination types. Furthermore, hallucination behaviors vary widely across tasks—such as question answering (zheng2023doeschatgptfallshort), summarization (cao2022hallucinated), and dialogue (das2022diving)—yet most prior work focuses narrowly on specific settings (mundler2023self), limiting scalability and robustness.
@@
Hallucinations in large language models (LLMs) have been extensively studied, focusing on their causes (pan2023risk; chen2024combating; kasai2024realtime; wang2023causal; lee2022factuality; yao2023editing), evaluation methodologies (lin2021truthfulqa; lee2022factuality; min2023factscore; li2023halueval), and behavioral analysis (zhao2023knowing; dong2024statistical; li2023benchmarking). Many studies have investigated ways to mitigate hallucinations through retrieval-augmented generation (peng2023check; varshney2023stitch; kang2023ever) and supervised fine-tuning (elaraby2023halo; razumovskaia2024dial; zhang2023r). To improve LLM reliability, researchers have explored prompting-based solutions. For example, si2022prompting proposed simple yet effective prompts that enhance GPT-3’s factual accuracy, while mitchell2022enhancing introduced a two-model framework in which one model generates responses and another evaluates their logical coherence. More recently, mundler2023self found that 17.7% of ChatGPT-generated sentences contain self-contradictions and proposed a three-step pipeline to detect and mitigate them without relying on external knowledge. Despite these advancements, hallucination detection and mitigation in LLMs remain challenging. Our method uses task-adaptive prompts in a zero-shot setting to guide LLMs in detecting and revising hallucinated content by eliciting and leveraging model reasoning.
@@
Prompting strategies have played a crucial role in enhancing the reasoning abilities of LLMs. Chain-of-Thought (CoT) prompting (wei2022chain) explicitly structures intermediate reasoning steps, significantly improving model performance on complex reasoning tasks. Building upon this, various enhancements have been proposed, including prompt ensembling (wang2022self; li2022advance; fu2022complexity), problem decomposition (zhou2022least; khot2022decomposed; dua2022successive), and structured planning methods (yao2022react; huang2022language; wang2023describe; liu2023llm+). To reduce manual effort and computational overhead, zero-shot CoT prompting (kojima2022large) was introduced, allowing LLMs to autonomously generate reasoning steps without the need for labeled exemplars. However, existing methods primarily focus on general reasoning tasks, and limited work has explored their applicability in addressing hallucinations within LLM-generated text.
@@
Contradictions within the same response (e.g., mutually exclusive statements) signal hallucination. Such contradictions appear in  17.7% of ChatGPT-generated sentences (mundler2023self).
@@
The core innovation of HalluClean lies in its structural reasoning mechanism for hallucination detection. While direct classification can identify obvious hallucinations, we find that step-by-step reasoning significantly improves detection accuracy, especially for subtle or complex hallucination cases. Our approach draws inspiration from cognitive science literature on human reasoning, which emphasizes the role of structured thinking in error detection and verification (kahneman2011thinking).
@@
We collect evaluation data from four established hallucination detection benchmarks: 1. HaluEval (li2023halueval): Covers hallucinated samples across three task types—question answering, knowledge-grounded dialogue, and text summarization. 2. UMWP (sunbenchmarking): Evaluates hallucination in math word problems (MWPs) by identifying questions with no or non-unique solutions. Such unanswerable questions are known to induce hallucinations in LLMs and are often used to test whether models can recognize ill-posed or unsolvable problems—similar to how educators gauge student understanding with trick questions. 3. ChatProtect (mundler2023self):Focuses on self-contradictory hallucinations, where a language model produces logically inconsistent statements within the same context. 4. HaluBench (ravi2024lynx):A domain-specific benchmark composed of hallucinated QA examples in the medical and financial domains, sourced from CovidQA, PubMedQA, and FinanceBench.
@@
The effectiveness of the proposed framework is evaluated along three dimensions: 1. Hallucination Reduction Rate: To measure the effectiveness of the revision step, we compute the hallucination reduction rate by comparing the number of hallucinations detected before and after revision. Specifically, we first identify hallucinations in the original outputs, then re-evaluate the revised outputs using GPT-4o-mini. 2. Revision Success Rate: To assess revision quality, we compute BERTScore between each revised output and its gold reference. A revision is considered acceptable if the BERTScore (DBLP:journals/corr/abs-1904-09675) exceeds 0.85(chosen to balance strict semantic fidelity and flexibility in surface expression). The revision success rate is the proportion of acceptable revisions among all hallucinations identified before revision. For multi-word problems (MWPs), revision quality evaluation is performed based on unanswerable reason categories, using exact match between predicted and gold labels. 3. Hallucination Detection: Since hallucination detection is formulated as a binary classification task, we evaluate its effectiveness using two standard metrics: F1 score and accuracy, computed against human-annotated gold labels from the original benchmarks. Accuracy reflects overall correctness, while F1 provides additional insight into the model’s balance between precision and recall, ensuring that both metrics are measured with respect to verified ground truth.
@@
To evaluate whether our method complements external knowledge, we test hallucination detection on the QA task from HalEval (li2023halueval) under two settings: vanilla (no external knowledge) and retrieval-augmented (with background knowledge).
@@ Conclusion
Limitations and Future Work
While HalluClean exhibits strong zero-shot generalization across diverse tasks, several avenues remain for improvement. One current limitation lies in its reliance on the reasoning capabilities of the underlying language model. Although our approach performs robustly with state-of-the-art LLMs, its effectiveness may vary when deployed in low-resource environments or with smaller models. This reflects a broader challenge inherent to prompt-based systems rather than a flaw unique to HalluClean.
To mitigate this, future work will explore incorporating lightweight verification modules or hybrid approaches that combine retrieval-augmented generation with structured reasoning. Additionally, fine-tuning smaller models with distilled reasoning traces from larger LLMs could offer a promising direction to improve performance while reducing computational costs.
Ethics Statement
In this study, we uses the APIs provided by OpenAI and DeepSeek, which are also strictly used for research purposes and in accordance with the terms and conditions stipulated by OpenAI and DeepSeek.
Experimental settings
We conduct experiments using five popular instruction-tuned LLMs: (1) GPT-3.5 (gpt-3.5-turbo-0125), (2) GPT-4o-mini, (3) LLaMA-3.1-70B-Instruct (dubey2024llama), (4) DeepSeek-V3 (liu2024deepseek), and (5) DeepSeek-R1. For GPT-series and DeepSeek models, we interact via API using default temperature settings. For Llama-3-70B, we use the BitsAndBytes library to run the model in 4-bit, following the official configuration recommended by the Meta Llama repository. All models are implemented in PyTorch 2.5.1 with CUDA 12.4 and executed on an NVIDIA A100-SXM4-80GB GPU.
Dataset Statistics
We evaluate our framework on four hallucination-related datasets, covering diverse tasks and domains. To ensure fairness, we maintain class balance when sampling examples for evaluation. The summary of the datasets, including their source, task type, and the number of used instances, is shown in Table 9. Here, pos refers to samples labeled as containing hallucinations.
Task NN $n_{01}$ $n_{10}$ $\chi^{2}$ $p$
QA 400 112 83 4.02 0.045
DA 400 84 71 0.93 0.335
SUM 400 94 40 20.96 $4.68\times 10^{-6}$
MWPs 400 121 33 49.15 $2.37\times 10^{-12}$
SC 300 89 20 42.42 $7.36\times 10^{-11}$
Overall 1900 500 247 85.01 $3.0\times 10^{-20}$
Table 8: McNemar tests for hallucination detection accuracy between GPT-3.5-turbo and HalluClean (Ours-GPT-3.5) across different tasks. $n_{01}$ and $n_{10}$ count instances where GPT-3.5 is incorrect while HalluClean is correct, and vice versa.
TASK Dataset Data Source #Instances(pos/neg)
Question Answering HalEval HotPotQA 400(200/200)
Dialogue HalEval OpenDiaKG 400(200/200)
Summarization HalEval CNN/DailyMail 400(200/200)
Math Word Problems UMWPs SVAMP/MultiArith/GSM8K/ASDiv 400(200/200)
Self Contradiction Chatprotect Wikipedia 300(150/150)
CovidQA Halubench scientific articles 200(100/100)
PubmedQA Halubench PubMed 200(100/100)
FinanceBench Halubench Companies’ public filings 200(100/100)
HalluQA(Chinese) HalluQA Chinese historical culture/customs/social phenomena 200(100/100)
CMHE-HD(Chinese) CMHE CMD/cMedQA2 200(100/100)
Table 9: Dataset statics
Significance test
we conduct paired significance tests comparing our framework with the GPT-3.5-turbo direct-ask baseline.
For hallucination detection, we treat accuracy as the primary metric and apply McNemar’s test to per-instance correctness. As shown in Table 8, HalluClean makes significantly more correct decisions than GPT-3.5-turbo on most tasks, and the overall improvement is highly significant ($\chi^{2}=85.01$, $p=3.0\times 10^{-20}$).
Demonstration Cases
Representative examples from five tasks—Question Answering (QA), Dialogue System (DA), Summarization (SUM), Math Word Problems (MWPs), and Self-contradictory hallucination (SC)—are shown in Figures 4 to 8, where our framework successfully detects and revises hallucinations.
Figure 4: An illustrative example of the question answering hallucination detection and revision. original hallucinated content, correct conclusions, and correct reasoning and revisions.
Figure 5: An illustrative example of the dialogue system hallucination detection and revision.original hallucinated content, correct conclusions, and correct reasoning and revisions.
Figure 6: An illustrative example of the summarization hallucination detection and revision. original hallucinated content, correct conclusions, and correct reasoning and revisions.
Figure 7: An illustrative example of the unanswerable math word problems detection and revision.original under-specified content, correct conclusions, and correct reasoning and revisions.
Figure 8: An illustrative example of the self-contradictory hallucination detection and revision. original self-contradiction content, correct conclusions, and correct reasoning and revisions.
Results on Module Applicability
To assess the generalizability of our framework components, we conduct module-level ablation experiments across different backbone LLMs. Specifically, we evaluate the performance of the detection module (Table 10) and the revision module (Table 11) when integrated with various base models.
Model QA DA SUM MWPs SC
F1 ACC F1 ACC F1 ACC F1 ACC F1 ACC
+ours 67.8% 66.5% 74.3% 69.3% 65.9% 69.2% 80.3% 81.5% 87.0% 87.0%
+ours 70.4% 69.0% 77.1% 72.5% 69.5% 69.3% 89.1% 89.8% 85.1% 74.0%
Llama-3-70B 44.7% 62.3% 66.4% 62.3% 32.4% 59.3% 83.4% 83.5% 65.8% 73.7%
+ours 70.6% 69.0% 74.0% 68.8% 46.5% 61.5% 85.6% 86.0% 80.8% 83.3%
+ours 71.5% 70.5% 77.1% 72.5% 62.9% 67.5% 89.1% 89.5% 76.1% 80.3%
Table 10: Detection module adaptability across backbone LLMs.
Model QA DA SUM MWP SC
R Q R Q R Q R Q R Q
GPT-3.5-turbo 20.5% 12.0% 57.5% 53.0% 14.5% 14.0% 42.0% 13.0% 30.7% 30.7%
+Ours 72.5% 25.5% 89.0% 83.0% 59.5% 59.0% 75.5% 34.0% 87.3% 87.3%
GPT-4o-mini 20.5% 12.0% 57.5% 51.5% 30.0% 14.5% 42.0% 19.0% 30.7% 30.7%
+Ours 70.5% 33.5% 89.0% 82.5% 59.5% 59.5% 75.5% 36.0% 87.3% 87.3%
Llama-3-70B 20.5% 14.0% 57.5% 52.5% 19.5% 14.5% 42.0% 18.5% 30.7% 30.7%
+Ours 70.5% 35.5% 89.0% 83.0% 59.5% 59.5% 75.5% 36.0% 87.3% 87.3%
DeepSeek-V3 20.5% 12.5% 57.5% 51.5% 41.0% 14.5% 42.0% 21.5% 30.7% 30.7%
+Ours 70.5% 36.0% 89.0% 83.0% 59.5% 59.5% 75.5% 36.5% 87.3% 87.3%
DeepSeek-R1 17.0% 12.5% 57.5% 50.0% 36.5% 14.0% 42.0% 25.0% 30.7% 30.7%
+Ours 70.5% 35.0% 89.0% 82.0% 59.5% 59.0% 75.5% 37.5% 87.3% 87.3%
Table 11: Revision module adaptability across backbone LLMs. R denotes the hallucination reduction rate after applying the revision module, and Q denotes the revision success rate, which reflects the quality of corrections.
Detailed Prompt Examples Used in Experiments
We present the exact prompts used in our experiments, including our structural-reasoning mechanism (Table 12), which encourage more structured and faithful reasoning, as well as baseline prompt strategies (Table 13) adapted from prior works for comparison.
Task Type Structural Reasoning Prompt
Question Answering Step-1 You are provided with a question and its corresponding answer. Your task is to determine whether the answer contains hallucinated content. Let’s understand the task and devise a plan to solve the task. [Task Input]
Step-2 Let’s carry out the plan and solve the task step by step. Show the reasoning process. [Task Input;Result from Step-1]
Step-3 Please conclude whether the answer contains hallucinated content with Yes or No. [Task Input;Result from Step-2]
Step-4 Given a question, its corresponding hallucinated answer, and an analysis explaining why the answer contains hallucinated content. Your task is to answer the question without introducing any hallucinations. [Task Input;Result from Step-2]
Dialogue Systems Step-1 You are provided with a dialogue history and its corresponding response. Your task is to determine whether the response contains hallucinated content. Let’s understand the task and devise a plan to solve the task. [Task Input]
Step-2 Let’s carry out the plan and solve the task step by step. Show the reasoning process. [Task Input;Result from Step-1]
Step-3 Please conclude whether the response contains hallucinated content with Yes or No. [Task Input;Result from Step-2]
Step-4 Given a dialogue history and its corresponding hallucinated response. Your task is to regenerate the response without introducing any hallucinations. [Task Input;Result from Step-2]
Summarization Step-1 You are provided with a document and its corresponding summary. Your task is to determine whether the summary contains hallucinated content. Let’s understand the task and devise a plan to solve the task. [Task Input]
Step-2 Let’s carry out the plan and solve the task step by step. Show the reasoning process. [Task Input;Result from Step-1]
Step-3 Please conclude whether the summary contains hallucinated content with Yes or No. [Task Input;Result from Step-2]
Step-4 Given a document,its corresponding hallucinated summary , and an analysis explaining why the summary contains hallucinated content. Your task is to regenerate the summary without introducing any hallucinations. [Task Input;Result from Step-2]
Math Word Problems Step-1 You are provided with a math word problem. Your task is to determine whether the problem is unanswerable. Let’s understand the task and devise a plan to solve the task. [Task Input]
Step-2 Let’s carry out the plan and solve the task step by step. Show the reasoning process. [Task Input;Result from Step-1]
Step-3 Please conclude whether the problem is unanswerable with Yes or No. [Task Input;Result from Step-2]
Step-4 Given a unanswerable math word problem and an analysis explaining why it is unanswerable. Your task is to revise the problem to make it answerable. [Task Input;Result from Step-2]
Self-contradiction You are given two texts. Your task is to determine whether the information in the two texts is contradictory. Let’s understand the task and devise a plan to solve the task. [Task Input]
Step-2 Let’s carry out the plan and solve the task step by step. Show the reasoning process. [Task Input;Result from Step-1]
Step-3 Please conclude whether the two texts are contradictory with Yes or No. [Task Input;Result from Step-2]
Step-4 Given Text 1, Text 2, and the analysis of the contradiction between them. Your task is to revise Text 2 to remove the contradiction, making it consistent with Text 1. [Task Input;Result from Step-2]
Table 12: Structured reasoning prompts for different NLP applications.
Prompt Strategy Prompt content
Direct Ask Please only conclude whether the text contains hallucinated content with Yes or No. [Question,Answer]
Step-by-step You are provided with a question and its corresponding answer. Does the answer contain hallucinated content? First, show your reasoning in a step-by-step fashion. Then conclude with yes or no.[Question,Answer]
SelfCheckGPT Is the sentence supported by the context above? Answer Yes or No.[Context:Question,Sentence:Answer]
Plan-and-Solve The Problem is to determine whether the text contains hallucinated content with Yes or No. Let’s first understand the problem and devise a plan to solve the problem.Then, let’s carry out the plan andsolve the problem step by step.[Question,Answer]
ChatProtect Step-1: You are provided with a question and its corresponding answer. Please explain whether the answer contains hallucinated content. Provide your explanation only.Step-2: Please conclude whether the answer contains hallucinated content with Yes or No. [Question,Answer]
Table 13: Baseline prompt strategies used for comparison with our detection method (QA task shown as example).
Error Analysis
We analyze the failure cases in hallucination detection, which can be broadly categorized into three types:
  • Misunderstanding of Language: The model fails to correctly interpret the question or the answer.
  • Lack of Background Knowledge: The model is unable to detect hallucinations when it lacks access to essential domain or factual knowledge required for verification.
  • Incorrect Reasoning: The model starts with a correct interpretation or premise but makes logical mistakes during the reasoning process, leading to incorrect conclusions.
In the following, we provide representative examples illustrating each of the three identified error types.
Figure 9: A typical case example of language misunderstanding, where the Red text highlights the part where misunderstanding of language causes the error.
Figure 10: A typical case of lacking background knowledge, with Red highlighting the error-causing segment due to missing related knowledge.
Figure 11: A typical case of incorrect reasoning, where the initially correct reason is highlighted in green, and the error in reasoning is marked in red.